Calculate the probabilities using geometric distribution Round to 3 decimal places.

a. The first success will occur on the 4 t h trial with p = 0.53 . P ( 4 ) =

b. The first success will occur on the 7 t h trial with p = 0.36 . P ( 7 ) =

c. The first success will occur on the 11 t h trial with p = 0.06 . P ( 11 ) =
Answer #1

If X denotes the number of trials to get first success, then X ~ Geometric(p)

The probability mass function of X is

PlX z) = p(1 _ p)-1, x = 0. 1. 2.3.

a) X ~ Geometric(p = 0.53)

The probability that the first success will occur on the 4th trial

P(X = 4) = 0.53(1-0.53)4-1 = 0.0550262

b) X ~ Geometric(p = 0.36)

The probability that the first success will occur on the 7th trial

P(X = 7) = 0.36(1-0.367-1-0.0247390

c)  X ~ Geometric(p = 0.06)

The probability that the first success will occur on the 11th trial

P(X = 11) = 0.06(1-0,06)11-1-00323169